(4x^3-3y^2)dx + 6xydy=0,f(-1)=0


1
Expert's answer
2021-10-10T09:33:25-0400
"6xyy'-3y^2=-4x^3"

"y'-\dfrac{1}{2x}y=-\dfrac{2}{3}x^2y^{-1}"

"v=y^{1-n}=y^{1-(-1)}=y^2"

"v'=2yy'"

"\dfrac{1}{2}v'-\dfrac{1}{2x}v=-\dfrac{2}{3}x^2"

"v'-\dfrac{1}{x}v=-\dfrac{4}{3}x^2"

The integrating factor


"\mu(x)=\dfrac{1}{x}"

"\dfrac{1}{x}v'-\dfrac{1}{x^2}v=-\dfrac{4}{3}x"

"d(\dfrac{v}{x})=-\dfrac{4}{3}xdx"

Integrate


"\int d(\dfrac{v}{x})=-\int \dfrac{4}{3}xdx"

"\dfrac{v}{x}=-\dfrac{2}{3}x^2+C"


"v=-\dfrac{2}{3}x^3+Cx"

"y^2=-\dfrac{2}{3}x^3+Cx"

"y=\pm\sqrt{-\dfrac{2}{3}x^3+Cx}"

Given "y(-1)=0"


"0=\pm\sqrt{-\dfrac{2}{3}(-1)^3+C(-1)}"


"C=\dfrac{2}{3}"

Answer:


"y(x)=-\sqrt{-\dfrac{2}{3}x^3+\dfrac{2}{3}x}"

or


"y(x)=\sqrt{-\dfrac{2}{3}x^3+\dfrac{2}{3}x}"